OpenStudy (anonymous):

What is the solution to the equation 3 ^(1 – 2x) = 2 ?

OpenStudy (anonymous):

I got -0.6309300955 as my answer but I don't think it's right...

OpenStudy (anonymous):

take log base three of both sides?

OpenStudy (anonymous):

I know it's like (1-2x)log3=log2

OpenStudy (anonymous):

I'm getting something closer to .2 on my calculator, but you should try solving with logarithms to get an exact value.

OpenStudy (anonymous):

Alright,thanks for the help!

OpenStudy (anonymous):

yeah, just change that above equation to \[1-2x = \log_{3} 2\] and you should get it.

OpenStudy (anonymous):

@sarahcoop3r - Your method is right so far. (1-2x)log(3)=log(2) Divide both sides by log(3) to get rid of it on the left side: \[1-2x=\frac{\log(3)}{\log(2)}\] Subtract 1 from both sides\[-2x=\frac{\log(3)}{\log(2)}-1\] And finally, divide by -2 to get x by itself: \[x=-\frac{1}{2}(\frac{\log(3)}{\log(2)}-1)\] You should be able to plug that into your calculator to get a final answer.